# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 3.1. Overview of the Complete Workflow
Based on Summary File 1 (SF1) tables, a database with the features and rows of the confidential Hundred-percent Detail File (HDF) is reconstructed, validated for agreement with the HDF and the confidential Census Edited File (CEF), linked to commercial databases and a quasi-identifier-only copy of the CEF (labeled $\text{CEF}_{atkr}$) to determine putative reidentifications, then reidentifications are confirmed by linkage to the full-feature CEF. SF1 is itself tabulated from the HDF using a processing sequence that begins with the Census Unedited File (CUF). See Figure 2.
Figure 1 provides a high-level overview of the databases that underlie the confidential and published versions of the 2010 Census and also summarizes our workflow using these databases. We begin by describing the internal Census Bureau databases from which the public SF1 was created. Next, we describe SF1 itself. We also describe the attributes and limitations of the commercial databases used in the analysis. We deliberately abstract from some of the complexity of these databases in order to focus on the features that are salient to our reconstruction and reidentification attacks.
@@
CUF x x x x y x x
CEF x x x x y x x
HDF x x z y x x
COMRCL x x x x y
${\text{CEF}_{atkr}}$ x x x x y
$\text{rHDF}_{b,t}$ x x z y x x
$\text{rHDF}_{b}$ x x x x x
Putative $\text{rHDF}_{b,t}$ x x x z y x x
Putative $\text{rHDF}_{b}$ x x x x x x
Confirmed $\text{rHDF}_{b,t}$ x x x z y x x
Confirmed $\text{rHDF}_{b}$ x x x x x x
MDG x x x z y g g
PRG x x x z y h h
MDF x x z y x x
${\text{rMDF}_{b,t}}$ x x z y x x
$\text{rSWAPLo}_{b,t}$ x x z y x x
$\text{rSWAPHi}_{b,t}$ x x z y x x
@@
The U.S. Constitution mandates a census of population conducted every ten years. Since the 1970 Census, these enumerations have collected primarily self-reported information on households and the individuals in those households. For the 2010 Census, the confidential respondent microdata were stored in several databases. For our purposes, describing the relevant feature sets and provenance requires starting with the 2010 Census Unedited File (CUF), which contains the raw census responses for all living quarters, unduplicated and deemed in-scope for the enumeration. The 2010 Census Edited File (CEF) constitutes the final, fully edited, permanent electronic record of these responses to the 2010 Census. The application of confidentiality protections and tabulation recodes to the CEF produces the Hundred-percent Detail File (HDF), which is used to create all published data products. The tabulation edits in HDF recode the residential location to the 2010 Census tabulation geography, create various age groupings, and create a variety of race and ethnicity groupings, all described below.
@@
As part of the internal confidentiality safeguards, the respondent’s name and address are stored on the CUF, not the CEF. Census data processing links the physical address to an identifier called the Master Address File Identifier (MAFID) that the Census Bureau’s Geography Division has determined to be a living quarter that existed on April 1, 2010, as either a housing unit or an occupied group quarters facility, and thus is in-scope for data collection in the 2010 Census. To facilitate research while safeguarding the name and address, the Census Bureau creates a crosswalk that relates the internal person-record identifier on the CEF to a person identifier called a Protected Identification Key (pik) using the production household data record-linkage system called the Person Identification Validation System (PVS). [^4] The PVS has evolved over time. The application of the production record linkage system was completed contemporaneously with the 2010 Census data processing using the 2010-vintage version of the PVS. For details on pik assignment, see [carra-2014-01, ]. While the feature sets for the full hierarchical CUF, CEF and HDF are much larger, the portions of the CUF that we use are shown in Table 1 in the row labeled “CUF.”
@@
The use of the Census Bureau’s production record-linkage system, and the selection of the 2010 Census vintage allowed us to do record linkage on name and address without having to design our own linkage system. We ensured that the same vintage of PVS was used for the commercial data we discuss later in this section. If the PVS recognized a person in the 2010 Census, it is extremely likely that the same vintage of the software would recognize the same person in the commercial data, thus assigning the same pik. All record linkage systems are subject to false positive and negative linkages. By employing the same production PVS system on all data used for this paper, we accept linkages based on piks with the error properties described in [carra-2014-02, ].
@@
The MAFID is further geocoded into the 2010 Census tabulation geography. In this paper, we distinguish two components of the 15-digit tabulation geography–census tract (11-digit concatenation of FIPS state, county equivalent, tract) and census block (15-digit concatenation of FIPS state, county equivalent, tract, block). [^5] FIPS stands for Federal Information Processing Standards, and refers to numeric and two-letter alphabetic codes defined in U.S. FIPS Publication 5-2. FIPS 5-2 was superseded by ANSI standard INCITS 38:2009. For details, see [ansifips, ]. Census blocks are a statistical definition of geography, not the commonly used “city block,” with complete coverage of the entire territory of the United States, and are the atoms in the Census Bureau’s geographic lattice that are used to build all other geographic tabulation summary levels, such as census tracts or county equivalents. Census blocks are defined in terms of territory, not population, and tessellate the entire United States. Some blocks may therefore be uninhabited (even underwater), others may have a very large population. See [censusblocks, ] for an overview. For more details on the person and geography attributes, see [sf1:2010:tech, ]. The HDF is formed from the CEF by applying the SDL methods described in Section 2.2, which are called “confidentiality edits” in the technical documentation [sf1:2010:tech]. Finally, the parts of the person records in CEF and HDF used in this paper have the feature sets shown in Table 1 in the appropriately labeled rows.
@@
The confidential databases share the same schema and feature sets: one column for the census block, one column each for the person identifier (pik), sex, age, and ethnicity (Hispanic or Lation/Not Hispanic or Latino), and six columns for the required race categories. [^6] For details on the person attributes, see [sf1:2010:tech, ]. For background on the definitions of the required race and ethnicity categories, see Statistical Policy Directive 15 [OMB:1997]. The category “some other race” is mandated by law, not statistical policy. Persons may self-declare multiple race categories; hence the binary race features are not mutually exclusive. In practice, most 2010 Census respondents only identified with a single race. The feature age is recorded in integer values. If a census response is missing, the process that creates the CEF performs edits and imputation, called “allocation” in the technical documents. There are no missing data in the CEF and, in particular, at least one of the six race categories must be selected. Excluding pik, which is standing in for name, we define all valid combinations of block, sex, age, race, and ethnicity as the feature space (sample space in statistics) for CEF and HDF, $\chi$. There are approximately $161\times 10^{9}$ (161 billion) such combinations which gives the cardinality $|\chi|$.777 Cardinality $|\chi|=161,109,592,812=6,207,027\times 2\times 103\times 63\times 2$, where $6,207,027$ is the number of inhabited blocks in the 2010 Census, 103 is the number of single-year age categories 0 to 99 plus grouped ages 100-104, 105-109, 110+ allowed in the published tables, and $63$ is the number of allowable race combinations. Note that, for technical reasons, when we implement the reconstructions of HDF, we modify the feature set to eliminate the age binning for ages 100+. Finally, note that we used data for the 50 states and the District of Columbia. [^8] The use of the term “state” in this document refers to all 51 state-equivalent political divisions. We excluded Puerto Rico because the 2010 vintage of the Census Bureau’s production name and address record linkage system did not work as well for this commonwealth.
@@
The most extensive and widely used 2010 Census data product is Summary File 1 [sf1:2010:tech]. Figure 2 illustrates the process of creating SF1 from the internal census databases. SF1 contains counts of persons, households, families within households, group quarters residents, and housing units tabulated at the census block, census tract, and county-equivalent geographic levels. SF1 also includes the tables released separately as the 2010 P.L. 94-171 Redistricting Data Summary File, which forms the basis for redistricting every legislative body in the United States and is normally released by March 31st of the year following the decennial census, several months before SF1.999The populations used for apportionment include a limited number of U.S. citizens and their families living abroad, known as the Federally Affiliated Overseas Population. These persons do not have records in the CEF, and their total for each state is added to the residential population for that state prior to apportionment, see Appendix G of [sf1:2010:tech, ]. The P.L. 94-171 Redistricting Data Summary File tables are renumbered in SF1 but are otherwise identical to the original release.
@@
The SF1 and other published tables are created by tabulating the HDF according to various combinations of geographic and demographic detail. All published data from the 2010 Census used the same geographic hierarchy. See Appendix A of [sf1:2010:tech, ] for more details. The census block is the most detailed geographic category. There were 11,078,297 blocks defined for 2010 Census publications of which 6,207,027 had nonzero populations. [^10] Whether a block was inhabited or not was published without any confidentiality edits in 2010. These census blocks aggregate into 73,057 defined census tracts of which 72,531 had nonzero populations. These census tracts, in turn, aggregate into 3,143 county equivalents, all of which had nonzero populations.
@@
P1 P6 P7 P8 P9 P10 P11 P12
P14 P12A P12B P12C P12D P12E P12F P12G
P12H P12I
@@
PCT12 PCT12A PCT12B PCT12C PCT12D PCT12E PCT12F PCT12G
PCT12H PCT12I PCT12J PCT12K PCT12L PCT12M PCT12N PCT12O
@@ 3.4. The Treatment of Age in Summary File 1
2010 SF1 tabulated age differently depending on the specific table and the level of geographic detail. At the tract level and above (e.g., Table PCT12), age was tabulated in single years from 0 to 99 years, then binned into the ranges 100-104 years, 105-109 years, and 110 years and over. At the block level in most tables (e.g., Table P12), age was binned into the following ranges: 0-4; 5-9; 10-14; 15-17; 18-19; 20; 21; 22-24; 25-29; 30-34; 35-39; 40-44; 45-49; 50-54; 55-59; 60-61; 62-64; 65-66; 67-69; 70-74; 75-79; 80-84; and 85+. Also at the block level, Tables P10 and P11 selected only persons age 18 and older. Finally, the block-level table P14 selected only individuals age 20 or younger and encoded age in single years. Combining the different age binning and universe selection rules applied at the census block level defines the most detailed age schema that these tables can support. That schema has 38 age groups: single year of age from 0 to 21, then: 22-24; 25-29; 30-34; 35-39; 40-44; 45-49; 50-54; 55-59; 60-61; 62-64; 65-66; 67-69; 70-74; 75-79; 80-84; and 85+. We use this 38-bin age schema (feature agebin) in our assessments of agreement of the reconstructed HDF with the HDF and CEF. We show in the reidentification experiments that the 38-bin age schema provides sufficient uniqueness for persons at the census block level to enable reconstruction-abetted reidentification. It is important to note that tables P12 and P14, when combined, give the exact number of males and females in each of the 38 age bins in each census block, so any remaining uncertainty is in the race and ethnicity distribution within each sex and age bin. For any reconstructed microdata based on at least the tables in Panel A of Table 2, there is only one possible reconstruction on the feature set {block, sex, agebin}; that is, the reconstruction is exact on that feature set. In our matching algorithms, we distinguish between matches based on exact age (feature age) and those based on binned age (feature agebin). In Section 4.5 we elaborate on the solution variability of our reconstructions.
@@ 3.5. Circa 2010 Commercial Databases
We created the commercial data (COMRCL) used for our reidentification experiments by combining data extracts originally purchased in support of the 2010 Census evaluations from four commercial providers between 2009 and 2011.121212The four commercial databases were provided by Experian Marketing Solutions Incorporated, Infogroup Incorporated, Targus Information Corporation, and VSGI LLC. The databases used are the same as in [rastogi2012, ] except that we excluded data provided by the Melissa Data Corporation, which contain address information but not sex and age data. The COMCRL data serve as the background knowledge of an attacker with low-quality information contemporaneous with the release of the 2010 SF1. While the database schema and the purposes for which these commercial databases were originally collected vary, they all share certain attributes. All have basic personal identifying information (PII) including names, addresses, sex, and birthdates. The vintage 2010 versions of these databases that we used did not include self-reported race and ethnicity data. [^13] Race and ethnicity data are modeled in some of the commercial databases [rastogi2012].
@@
(×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$)

### v2

@@ 3.1. Overview of the complete workflow
Based on Summary File 1 (SF1) tables, a database with the features and rows of the confidential Hundred-percent Detail File (HDF) is reconstructed, validated for agreement with the HDF and the confidential Census Edited File (CEF), linked to commercial databases and a quasi-identifier-only copy of the CEF (labeled $\text{CEF}_{atkr}$) to determine putative reidentifications, then reidentifications are confirmed by linkage to the full-feature CEF. SF1 is itself tabulated from the HDF using a processing sequence that begins with the Census Unedited File (CUF), see Figure 2.
Figure 1 provides a high-level overview of the databases that underlie the confidential and published versions of the 2010 Census and also summarizes our workflow using these databases. We begin by describing the internal Census Bureau databases from which the public SF1 was created. Next, we describe SF1 itself. We also describe the attributes and limitations of the commercial databases used in the analysis. We deliberately abstract from some of the complexity of these databases in order to focus on the features that are salient to our reconstruction and reidentification attacks. Table 1 lists all data used or produced by the workflow in Figure 1.
@@
CUF (Census Unedited File) x x x x y x x
CEF (Census Edited File) x x x x y x x
HDF (Hundred-Percent Detail File) x x z y x x
COMRCL (Commercial Attack File) x x x x y
${\text{CEF}_{atkr}}$ (CEF, Attack Columns) x x x x y
$\text{rHDF}_{b,t}$ (Single-Year-Age Reconstruction) x x z y x x
$\text{rHDF}_{b}$ (Binned-Age Reconstruction) x x x x x
Putative $\text{rHDF}_{b,t}$ (With All Links) x x x z y x x
Putative $\text{rHDF}_{b}$ (With All Links) x x x x x x
Confirmed $\text{rHDF}_{b,t}$ (Confirmed Links) x x x z y x x
Confirmed $\text{rHDF}_{b}$ (Confirmed Links) x x x x x x
MDG (Modal Guess Attacker) x x x z y g g
PRG (Proportional Guess Attacker) x x x z y h h
MDF (2020 DAS Output Microdata) x x z y x x
${\text{rMDF}_{b,t}}$ (MDF Reconstruction) x x z y x x
$\text{rSWAPLo}_{b,t}$ (Low-Swap Reconstruction) x x z y x x
$\text{rSWAPHi}_{b,t}$ (High-Swap Reconstruction) x x z y x x
@@
The U.S. Constitution mandates a census of population conducted every ten years. Since the 1970 Census, these enumerations have collected primarily self-reported information on households and the individuals in those households. For the 2010 Census, the confidential respondent microdata were stored in several databases. For our purposes, describing the relevant feature sets and provenance requires starting with the 2010 Census Unedited File (CUF), which contains the raw census responses for all living quarters, unduplicated and deemed in-scope for the enumeration. The 2010 Census Edited File (CEF) constitutes the final, fully edited (e.g., for logical consistency between the ages of parents and their biological children), permanent electronic record of these responses to the 2010 Census. The application of confidentiality protections and tabulation recodes to the CEF produces the Hundred-percent Detail File (HDF), which is used to create all published data products. The tabulation edits in HDF recode the residential location to the 2010 Census tabulation geography, create various age groupings, and create a variety of race and ethnicity groupings, all described below.
@@
As part of the internal confidentiality safeguards, the respondent’s name and address are stored on the CUF, not the CEF. Census data processing links the physical address to an identifier called the Master Address File Identifier (MAFID) that the Census Bureau’s Geography Division has determined to be a living quarter that existed on April 1, 2010, as either a housing unit or an occupied group quarters facility, and thus is in-scope for data collection in the 2010 Census. To facilitate research while safeguarding the name and address, the Census Bureau creates a crosswalk that relates the internal person-record identifier on the CEF to a person identifier called a Protected Identification Key (pik) using the production household data record-linkage system called the Person Identification Validation System (PVS). [^5] The PVS has evolved over time. The application of the production record-linkage system was completed contemporaneously with the 2010 Census data processing using the 2010-vintage version of the PVS. For details on pik assignment, see carra-2014-01. While the feature sets for the full hierarchical CUF, CEF and HDF are much larger, the portions of the CUF that we use are shown in Table 1 in the row labeled “CUF.”
@@
The use of the Census Bureau’s production record-linkage system, and the selection of the 2010 Census vintage allowed us to do record linkage on name and address without having to design our own linkage system. We ensured that the same vintage of PVS was used for the commercial data we discuss later in this section. If the PVS recognized a person in the 2010 Census, it is extremely likely that the same vintage of the software would recognize the same person in the commercial data, thus assigning the same pik. All record-linkage systems are subject to false positive and negative linkages. By employing the same production PVS system on all data used for this paper, we accept linkages based on piks with the error properties described in carra-2014-02.
@@
The MAFID is further geocoded into the 2010 Census tabulation geography. In this paper, we distinguish two components of the 15-digit tabulation geography–census tract (11-digit concatenation of FIPS state, county equivalent, tract) and census block (15-digit concatenation of FIPS state, county equivalent, tract, block). [^6] FIPS stands for Federal Information Processing Standards, and refers to numeric and two-letter alphabetic codes defined in U.S. FIPS Publication 5-2. FIPS 5-2 was superseded by ANSI standard INCITS 38:2009. For details, see ansifips. Census blocks are a statistical definition of geography, not the commonly used “city block,” with complete coverage of the entire territory of the United States, and are the atoms in the Census Bureau’s geographic lattice that are used to build all other geographic tabulation summary levels, such as census tracts or county equivalents. Census blocks are defined in terms of territory, not population, and tessellate the entire United States. Some blocks may therefore be uninhabited (even underwater), others may have a very large population. See censusblocks for an overview. For more details on the person and geography attributes, see sf1:2010:tech. The HDF is formed from the CEF by applying the SDL methods described in Section 2.3, which are called “confidentiality edits” in the technical documentation (sf1:2010:tech). Finally, the parts of the person records in CEF and HDF used in this paper have the feature sets shown in Table 1 in the appropriately labeled rows.
@@
The confidential databases share the same schema and feature sets: one column for the census block, one column each for the person identifier (pik), sex, age, and ethnicity (Hispanic or Latino/Not Hispanic or Latino), and six columns for the required race categories. [^7] For details on the person attributes, see sf1:2010:tech. For background on the definitions of the required race and ethnicity categories, see Statistical Policy Directive 15 (OMB:1997). The category “some other race” is mandated by law, not statistical policy (13 U.S. Code § 5 as amended in Pub. L. 111–117, div. B, title I, Dec. 16, 2009). Persons may self-declare multiple race categories; hence the binary race features are not mutually exclusive. In practice, most 2010 Census respondents only identified with a single race. [^8] Throughout this paper, the notation {race, ethnicity} refers to the the subset of features defined by the 63-category race and 2-category ethnicity features. The notation $race\times ethnicity$ distribution refers to the $63\times 2$ table using the 63-category race feature crossed with the 2-category ethnicity feature. The coding schemes are derived from the definitions used to create lines P8 and P9 in Table 2, Panel A. The feature age is recorded in integer values. If a census response is missing, the process that creates the CEF performs edits and imputation, called “allocation” in the technical documents. There are no missing data in the CEF and, in particular, at least one of the six race categories must be selected. Excluding pik, which is standing in for name, we define all valid combinations of block, sex, age, race, and ethnicity as the feature space (sample space in statistics) for CEF and HDF, $\chi$. There are approximately $161\times 10^{9}$ (161 billion) such combinations which gives the cardinality $|\chi|$.999 Cardinality $|\chi|=161,109,592,812=6,207,027\times 2\times 103\times 63\times 2$, where $6,207,027$ is the number of inhabited blocks in the 2010 Census, 103 is the number of single-year age categories 0 to 99 plus grouped ages 100-104, 105-109, 110+ allowed in the published tables, and $63$ is the number of allowable race combinations. Note that, for technical reasons, when we implement the reconstructions of HDF, we modify the feature set to eliminate the age binning for ages 100+. Finally, note that we used data for the 50 states and the District of Columbia. [^10] The use of the term “state” in this document refers to all 51 state-equivalent political divisions. We excluded Puerto Rico because the 2010 vintage of the Census Bureau’s production name and address record-linkage system did not work as well for this commonwealth.
@@
The most extensive and widely used 2010 Census data product is Summary File 1 (sf1:2010:tech). Figure 2 illustrates the process of creating SF1 from the internal census databases. SF1 contains counts of persons, households, families within households, group quarters residents, and housing units tabulated at the census block, census tract, and county-equivalent geographic levels. SF1 also includes the tables released separately as the 2010 Redistricting Data (Public Law 94-171) Summary File, which forms the basis for redistricting every legislative body in the United States and is normally released by March 31st of the year following the decennial census, several months before SF1.111111The populations used for apportionment include a limited number of U.S. citizens and their families living abroad, known as the Federally Affiliated Overseas Population. These persons do not have records in the CEF, and their total for each state is added to the residential population for that state prior to apportionment, see Appendix G of sf1:2010:tech. The Redistricting Data (Public Law 94-171) Summary File tables are renumbered in SF1 but are otherwise identical to the original release.
@@
The SF1 and other published tables are created by tabulating the HDF according to various combinations of geographic and demographic detail. All published data from the 2010 Census used the same geographic hierarchy. See Appendix A of sf1:2010:tech for more details. The census block is the most detailed geographic category. There were 11,078,297 blocks defined for 2010 Census publications of which 6,207,027 had nonzero populations. [^12] Whether a block was inhabited or not was published without any confidentiality edits in 2010. These census blocks aggregate into 73,057 defined census tracts of which 72,531 had nonzero populations. These census tracts, in turn, aggregate into 3,143 county equivalents, all of which had nonzero populations.
@@
Table Name Census Block Table Title
P1 Total Population
P6 Race (6 categories, alone or in combination)
P7 Hispanic or Latino Origin by Race
P8 Race (63 categories)
P9 Hispanic or Latino, and not Hispanic or Latino by Race
P10 Race for the Population 18 Years and Over
P11 Hispanic or Latino, and not Hispanic or Latino by Race
for the Population 18 Years and Over
P12 Sex by Age
P12A-I Sex by Age (iterated by Race and Ethnicity)
P14 Sex by Age for the Population Under 20 Years
@@
Table Name Census Tract Table Title
PCT12 Sex by Age
PCT12A-O Sex by Age (iterated by Race and Ethnicity)
@@ 3.4. The treatment of age in Summary File 1
2010 SF1 tabulated age differently depending on the specific table and the level of geographic detail. At the tract level and above (e.g., Table PCT12), age was tabulated in single years from 0 to 99 years, then binned into the ranges 100-104 years, 105-109 years, and 110 years and over. At the block level in most tables (e.g., Table P12), age was binned into the following ranges: 0-4; 5-9; 10-14; 15-17; 18-19; 20; 21; 22-24; 25-29; 30-34; 35-39; 40-44; 45-49; 50-54; 55-59; 60-61; 62-64; 65-66; 67-69; 70-74; 75-79; 80-84; and 85+. Also at the block level, Tables P10 and P11 selected only persons age 18 and older. Finally, the block-level table P14 selected only individuals age 20 or younger and encoded age in single years. Combining the different age binning and universe selection rules applied at the census block level defines the most detailed age schema that these tables can support. That schema has 38 age groups: single year of age from 0 to 21, then: 22-24; 25-29; 30-34; 35-39; 40-44; 45-49; 50-54; 55-59; 60-61; 62-64; 65-66; 67-69; 70-74; 75-79; 80-84; and 85+. We use this 38-bin age schema (feature agebin) in our assessments of agreement of the reconstructed HDF with the HDF and CEF. We show in the reidentification experiments that the 38-bin age schema provides sufficient uniqueness for persons at the census block level to enable reconstruction-abetted reidentification. It is important to note that tables P12 and P14, when combined, give the exact number of males and females in each of the 38 age bins in each census block, so any remaining uncertainty is in the race and ethnicity distribution within each sex and age bin. For any reconstructed microdata based on at least the tables in Panel A of Table 2, there is only one possible reconstruction on the feature set {block, sex, agebin}; that is, the reconstruction is exact on that feature set. In our matching algorithms, we distinguish between matches based on exact age (feature age) and those based on binned age (feature agebin). In Section 4.5 we present the formal mathematics for the solution variability of our reconstructions.
@@ 3.5. Circa 2010 commercial databases
We created the commercial data (COMRCL) used for our reidentification experiments by combining data extracts originally purchased in support of the 2010 Census evaluations from four commercial providers between 2009 and 2011.141414The four commercial databases were provided by Experian Marketing Solutions Incorporated, Infogroup Incorporated, Targus Information Corporation, and VSGI LLC. The databases used are the same as in rastogi2012 except that we excluded data provided by the Melissa Data Corporation, which contain address information but not sex and age data. The COMRCL data serve as the background knowledge of an attacker with lower-quality information contemporaneous with the release of the 2010 SF1. While the database schema and the purposes for which these commercial databases were originally collected vary, they all share certain attributes. All have basic personal identifying information (PII) including names, addresses, sex, and birth dates. The vintage 2010 versions of these databases that we used did not include self-reported race and ethnicity data. [^15] Race and ethnicity data are modeled in some of the commercial databases (rastogi2012).
@@
(×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$)